# Depth-Width Tradeoffs in Algorithmic Reasoning of Graph Tasks with Transformers

### v1

@@ 4.1 Improving the sub-linear lower bounds
Previous works on the connection between the MPC model and transformers have shown conditional lower bounds for solving certain tasks. One such lower bound include the one-cycle versus two-cycles problem. Namely, given a graph with $n$ nodes, distinguish whether it is one cycle of length $n$ or two cycles of length $\frac{n}{2}$. Conjecture 13 from Sanford et al. (2024a) (see also (Ghaffari et al., 2019)) says that the MPC model with $n^{1-\epsilon}$ memory per machine for any $\epsilon\in(0,1)$ cannot distinguish between the two cases, unless the number of MPC rounds is $\Omega(\log(n))$. This also suggests that transformers with an embedding dimension of $O(n^{1-\epsilon})$ cannot solve this task. We will first show that having a transformer with linear embedding dimension can solve this task:
@@
The proof intuition is that since the graph contains only $n$ edges, we can stack all of them into a single token. After doing so, we can use the MLP to solve it. This result shows that these kind of lower bounds for transformers are brittle in the sense that increasing the embedding dimension even slightly breaks them. With that said, this result is not surprising, since the graph in this task is very sparse.
The connectivity problem for general graphs cannot be solved in this manner.
The reason is that for graphs with $n$ nodes there are possibly $\Omega(n^{2})$ edges, and thus it is not possible to compress the entire graph into a single token with linear embedding dimension. However, the connectivity problem can still be solved by increasing the embedding dimension by polylog terms, as we explain below.
Ahn et al. (2012) uses linear sketching to solve the connectivity problem on general graphs with $O(n\log^{3}(n))$ total memory. The idea is to use linear sketching, which is a linear projection of the adjacency rows into vectors of dimension $O(\text{poly}\log(n))$. Although this is a lossy compression, it still allows to solve the connectivity problem. As a consequence, we can use a similar construction as in Theorem 4.1 where we first apply the sketching transformation to each token (i.e. row of the adjacency), then embed all the tokens, into a single token and use the MLP to solve the problem using the algorithm from Ahn et al. (2012) (Theorem 3.1).
Our next results will include problems that cannot be solved by simply compressing all the information of the graph into a single token, and would require a more intricate use of the self-attention layers.
4.2 The power of a linear embedding dimension
We have shown that transformers with a linear embedding dimension can improve certain lower bounds that apply for sub-linear dimensions. In this section we will show another lower bound on the task of detecting small cycles. This lower bound is stronger than the one for the one-cycle versus two-cycles in the sense that the depth-embedding dimension trade-off is even more strict.
Consider the task of detecting $2$-cycles in a directed graph. Namely, detecting whether there are nodes $u$ and $v$ with two edges, one directed from $u$ to $v$ and another directed from $v$ to $u$. We can prove the following lower bounds on this task:
@@ 4.2 The power of a linear embedding dimension
Let $T$ be a transformer with embedding dimension $m$ depth $L$, bit-precision $p$ and $H$ attention heads in each layer. Also, assume that the input graphs to $T$ are embedded such that each token is equal to a row of the adjacency matrix. Then, if $T$ can detect $2$-cycles on directed graphs we must have that:
  1. 1.
    If $T$ has residual connections then $mpHL=\Omega(n)$.
  2. 2.
    If $T$ doesn’t have residual connections then $mpH=\Omega(n)$.
The proof uses a communication complexity argument, and specifically a reduction to set disjointness problem. For a formal definition see Appendix B. This lower bound is stronger than the lower bounds in Section 4.1 in the sense that they cannot be improved by having logarithmic depth.
For example, assume that the embedding dimension is $m=O(n^{1-\epsilon})$ for some $\epsilon\in(0,1)$, and that $p,H=O(\log(n))$ (which is often the case in practice). Then, transformers with residual connection require a depth of $\Omega(n^{\epsilon})$ to solve this task, which is beyond logarithmic, while transformer without residual connections cannot solve it. This lower bound is also unconditional, compared to the lower bouds from (Sanford et al., 2024a) which are conditional on the hardness of the one-cycle versus two-cycles problem (Conjecture 13 there). The caveat of this lower bound is that it relies on having the input graph specifically embedded as rows of the adjacency matrix.
We now turn to show an upper bound for this task. We will show an even more general claim than detecting $2$-cycles. Namely, that depth $L$ transformers with $\Omega(n)$ embedding dimension can calculate the $L$-th power of an adjacency matrix of graphs. In particular, the trace of $A^{L}$ provides the number of cycles of size $L$ in the graph (multiplied by an appropriate constant). Thus, a $2$-layer transformer with linear embedding dimension can already solve the directed $2$-cycle problem.
@@
The constructive proof of this theorem carefully selects key and query parameters to ensure that the output of the softmax matrix equals $A$. This enables an inductive argument that encodes $A^{\ell}$ in the value matrix of the $\ell$th layer, in order to compute $A^{\ell+1}$.
The above theorem shows that having a transformer with linear embedding dimension in the graph size is very strong in terms of expressivity. Namely, $A^{L}_{i,j}$ counts the number of walks of length $L$ form node $i$ to node $j$. This also allows to determine whether a graph is connected, by checking whether $A^{n}$ doesn’t contain any zero entries. Although, there are other algorithm (e.g. those presented in Section 4.1) that can solve the connectivity task more efficiently.
4.3 Sublinear embedding dimension for bounded-degree graphs
The results of the previous section establish that for worst-case graph instance, transformers with node-adjacency tokenizations require a linear embedding dimension to solve simple graph reasoning tasks, such as 2-cycle detection. This fact is perhaps unsurprising because each node-adjacency tokenization input is a length-$n$ boolean vector, which must be compressed in a lossy manner if the embedding dimension $m$ and bit precision $p$ satisfy $mp=o(n)$. It is natural to ask whether such results apply to sparse graphs as well, such as graphs with bounded degree.
Here, we show that requisite embedding dimension to detect 2-cycles scales linearly with the degree of the graph.
@@ 4.3 Sublinear embedding dimension for bounded-degree graphs
For any $n\in\mathbb{N}$ and $d\leq n$, there exists a single-layer transformer with embedding dimension $O(d\log n)$ that detects 2-cycles in any graph with node degree at most $d$. This embedding dimensional is optimal up to logarithmic factors.
The proof reduces the dimensionality of the input adjacency matrix by incorporating vector embeddings from (Sanford et al., 2024c) into the key, query, and value matrices to produce a “sparse attention unit,” whose activations are large when a respective cycle exists.
5 Super-linear embedding dimension
In the previous section we have shown the power of having a transformer with an embedding dimension that is linear in the graph size. In this section we study transformers with a super-linear dimension.
Permitting the embedding dimension to $\Omega(n^{2})$ trivializes the question of expressive power in our setting.
In this regime, the entire graph can be encoded into a single token, which offsets any algorithmic demands to the computationally-unbounded MLP.
As a result, we restrict this section’s focus to embedding dimensions that are super-linear, but sub-quadratic, in the number of nodes.
5.1 Detecting and counting subgraphs
A very basic graph task is to detect or count a given subgraph. This task is important in several fields including biology, organic chemistry, graph kernels and more (See (Jin et al., 2020; Jiang et al., 2010; Pope et al., 2018; Shervashidze et al., 2009) and the discussion in Chen et al. (2020)). This task was also studied in the context of GNNs in Chen et al. (2020), where it was shown that $k$-IGN can detect subgraphs of size $k$. The caveat of this result is that $k$-IGN require $k$-tensor to construct, which is practically infeasible even for small $k$.
Here we study the ability of transformer to detect and count subgraphs. Our main result is the following:
@@ 5.1 Detecting and counting subgraphs
Let $k,n\in\mathbb{N}$, and let $G^{\prime}$ be a graph with $k$ nodes. There exists a transformer with $O(1)$ self-attention layers and embedding dimension $O\left(n^{2-1/k}\right)$ such that for any graph $G$ of size $n$,
counts the number of occurrences of $G^{\prime}$ as a subgraph of $G$.
The crux of the proof is to implement the seminal ”tri-tri-again“ algorithm (Dolev et al., 2012) using transformers. The general idea is that given a graph with $n$ nodes, we split the nodes into $n^{1/k}$ disjoint sets, each containing $n^{1-1/k}$ nodes, where $k$ is the size of the subgraph that should be detected or counted. For each possible combination of $k$ such sets we use an MLP to count the given subgraph in it. There are $\binom{n^{1/k}}{k}\leq n^{k\cdot 1/k}=n$ such combinations of sets, each one contains at most $n^{2-2/k}$ edges. Thus, each token with a large enough embedding dimension can simulate one combination of subsets, cumulating all the relevant edges. Note that subgraph detection is a sub-task of counting, whether the number of occurrences is larger than 00.
Theorem 5.1 can be compared to Theorem 23 in Sanford et al. (2024a) that provides a construction for counting $k$-cliques using transformers with sub-linear memory and additional blank tokens. There, it was shown that it is possible to count $k$-cliques with a transformer of depth $O(\log\log(n))$, however the number of blank tokens is $O(n^{k-1})$ in the worse case. By blank tokens we mean additional empty tokens that are appended to the input and used for scratch space as defined in (Sanford et al., 2024a). In our result, we require a depth of $O(1)$, and the total number of tokens is $n$, while the embedding dimension is super-linear (but sub-quadratic). Thus, our solution has better memory usage, since the increase in width is only polynomial, but the number of tokens is $n$ instead of $O(n^{k})$.
@@
5.2 Limitations of sub-quadratic memory
We now turn to the boundary of the possible embedding dimension for which these kinds of questions are interesting. A natural question is whether there is a width that suffices for solving all graph problems. Intuitively, this width should be n2 as this is what is required for storing the entire graph. Here we show that this is indeed an upper bound on needed depth for an arbitrary problem, and that this bound is tight for a particular problem: the Eulerian cycle verification problem.
Given a graph and a list of “path fragments,” where each fragment consists of a pair of subsequent edges in a path, the goal of the Eulerian cycle verification problem is to determine whether the properly ordered fragments comprise an Eulerian cycle. Recall that a Eulerian cycle is a cycle over the entire graph that uses each edge exactly once.
We show that this problem on multigraphs with self loops cannot be solved (conditioned on the $1$ vs. $2$ cycle conjecture) with a transformers unless its embedding dimensions is quadratic in $n$ or its depth is $\log(n)$.
Theorem 5.3.
@@ 6 Experiments
In this section, we present experiments to support our claims and theoretical results.  [^1] Code is provided in the Supplementary Material. In our experiments, we used a standard transformer architecture using Pytorch’s transformer encoder layers (Paszke et al., 2019). Specifically, each layer is composed of Multi-Head Self-Attention, Feedforward Neural Network, Layer Normalization and Residual Connections. More experimental details are provided in Appendix E.
For all experiments in Section 6.1 and Section 6.2, we used the adjacency rows tokenization as described in Section 5.1. Details of the implementation are described in Appendix D. We considered the tasks of connectivity, triangles count, and 4-cycle count. For the counting tasks, we used the substructure counting dataset from Chen et al. (2020), where each graph was labeled with the number of pre-defined substructures it contains, as a graph regression task. For the connectivity tasks, we generated synthetic graphs, and the label indicates whether the graph is connected or not. All the datasets information is described in detail in  Appendix E.
@@ 6.1 Empirical Trade-Offs Between Width and Depth
We trained a transformer with a fixed amount of 100k parameters split between varying depth and width. We examine how the running time, loss, and generalization depend on the width and depth. We examine the following pairs of (depth, width): $(1,125),(2,89),(4,63),(8,45),(10,40)$. We train each model for $100$ epochs and examine the following: the total training time, the total inference time, the training loss and test performance (accuracy for classification and MAE for regression). We repeat this experiment with graph sizes 50 and 100. We report the averages over $3$ runs with random seeds. The hyper-parameters we tuned are provided in Appendix E.
The results for the connectivity task over $100$ nodes are presented in Figure 2. Additional results for the counting tasks and different graph sizes, present the same trends, and are provided in Appendix D due to space limitations. As shown in Figure 2 and Figure 2, the training loss and accuracy remain consistent across all depth and width configurations. However, Figure 2 reveals that shallow and wide transformers significantly reduce the total training and inference time compared to their deeper and narrower counterparts. This may be due to the ability of GPUs to parallelize the computations across the width of the same layer, but not across different layers.
6.2 Critical Width
In this subsection, we show that the transformer width required to fit the substructure counting tasks increases roughly linearly with the number of nodes, as argued in  Section 5.1.
The experiments considered graphs with increasing numbers of nodes, ranging from 50 to 400 in increments of 50, and transformer widths varying from 100 to 800 in increments of 100. For each combination of graph size and transformer width, we determined the critical width at which the model failed to fit the data. The critical width is defined as the width where the training loss plateaued at more than 0.05. To determine the critical width, we conducted a grid search over each combination of graph size and model width, and selected the model that fitted the data best.
The hyper-parameters we considered are provided in Appendix E.
To isolate the effect of width, we used a transformer model with one layer. We used two attention heads to ensure there exists a width for which the model can fit the data for all the evaluated graph sizes. The results for the 4-Cycle count task are presented in Figure 3. Due to space limitations, the results for the triangle count task are deferred to Appendix D. This figure shows that the critical width increases roughly linearly with the graph size.

### v2

@@ 4 The Expressive Power of a Linear Embedding Dimension
Previous works on the connection between the MPC model and transformers have shown conditional lower bounds for solving certain tasks. One such lower bound includes the $1$ vs. $2$ cycle problem. Namely, determine whether an $n$-node graph is one cycle of length $n$ or two cycles of length $\frac{n}{2}$. Conjecture 13 from Sanford et al. (2024a) (see also Ghaffari et al. (2019)) states that any MPC protocol with $n^{1-\epsilon}$ memory per machine for any $\epsilon\in(0,1)$ cannot distinguish between the two cases, unless the number of MPC rounds is $\Omega(\log n)$. This condition implies that transformers with an embedding dimension of $O(n^{1-\epsilon})$ cannot solve this task. For a formal definition, see section˜B.1. We demonstrate the tightness of this bound by showing that a transformer with linear embedding dimension can solve this task:
@@
The proof intuition is that the graph contains only $n$ edges, we can stack all of them into a single token. After doing so, we can use the MLP to solve it. This demonstrates the brittleness of communication-based transformer lower bounds, since even a slight increase in the embedding dimension breaks them. With that said, this result is not surprising, since the graph in this task is very sparse.
The connectivity problem for general graphs cannot be solved in this manner. The reason is that for graphs with $n$ nodes there are possibly $\Omega(n^{2})$ edges, and thus it is not possible to compress the entire graph into a single token with linear embedding dimension. However, the connectivity problem can still be solved by increasing the embedding dimension by polylog terms, as we explain below.
Ahn et al. (2012) uses linear sketching to solve the connectivity problem on general graphs with $O(n\log^{3}(n))$ total memory. The idea is to use linear sketching, which is a linear projection of the adjacency rows into vectors of dimension $O(\text{poly}\log(n))$. Although this is a lossy compression, it still allows to solve the connectivity problem. As a consequence, we can use a similar construction as in Theorem˜4.1 where we first apply the sketching transformation to each token (i.e. row of the adjacency), then embed all the tokens, into a single token and use the MLP to solve the problem using the algorithm from Ahn et al. (2012) (Theorem 3.1). Note that this construction works only in high probability rather than deterministically, since the linear sketching requires using a random matrix, which compresses successfully the adjacency matrix only with some high probability. Thus, we don’t present it here formally, but rather as a proof sketch.
Our next results include problems that cannot be solved by simply compressing all the information of the graph into a single token, and require a more intricate use of the self-attention layers.
4.2 Linear width is necessary and sufficient for $2$-cycle detection
This section shows that for the problem of cycle detection, linear width is sufficient for solving the task with fixed-depth, and is also necessary. We consider the 2-cycle detection problem for directed graphs. The task is to find nodes $u$ and $v$ such that the edges $(u,v)$ and $(v,u)$ exist. We begin with the lower bound for this case.
@@ 4.2 Linear width is necessary and sufficient for $2$-cycle detection
Let $T$ be a transformer with embedding dimension $m$ depth $L$, bit-precision $p$ and $H$ attention heads in each layer. Also, assume that the input graphs to $T$ are embedded such that each token is equal to a row of the adjacency matrix. Then, if $T$ can detect $2$-cycles on directed graphs, the following must hold: (1) If $T$ has residual connections then $mpHL=\Omega(n)$; or (2) If $T$ doesn’t have residual connections then $mpH=\Omega(n)$.
The proof uses a communication complexity argument, and specifically a reduction to the set disjointness problem. For a formal definition see Appendix˜B. This lower bound is stronger than the lower bounds in Section˜4.1 in the sense that they cannot be circumvented by logarithmic depth. For example, assume that the embedding dimension is $m=O(n^{1-\epsilon})$ for some $\epsilon\in(0,1)$, and that $p,H=O(\log(n))$ (which is often the case in practice). Then, transformers with residual connection require a depth of $\Omega(n^{\epsilon})$ to solve this task, which is beyond logarithmic, while transformers without residual connections cannot solve it. This lower bound is also unconditional, compared to the lower bounds from Sanford et al. (2024a) which are conditional on the hardness of the $1$ vs. $2$ cycle problem (Conjecture 13 therein). The caveat of this lower bound is that it relies on having the input graph specifically embedded as rows of the adjacency matrix.
We now turn to show an upper bound for this task. We will show an even more general claim than detecting $2$-cycles. Namely, that depth $L$ transformers with $\Omega(n)$ embedding dimension can calculate the $L$-th power of an adjacency matrix of graphs. In particular, the trace of $A^{L}$ equals the number of cycles of size $L$ in the graph (multiplied by an appropriate constant). Thus, a $2$-layer transformer with linear embedding dimension can already solve the directed $2$-cycle problem.
@@
[…]
The above theorem shows that having a transformer with embedding dimension proportional to the graph size is representationally powerful. Namely, $A^{L}_{i,j}$ counts the number of walks of length $L$ form node $i$ to node $j$. This also allows to determine whether a graph is connected, by checking whether $A^{n}$ doesn’t contain any zero entries. Although, there are other algorithm (e.g. those presented in Section˜4.1) that can solve the connectivity task more efficiently.
4.3 A sublinear width solution for cycle detection in bounded degree graphs
The results of the previous section establish that for worst-case graph instance, transformers with node-adjacency tokenizations require a linear embedding dimension to solve simple graph reasoning tasks, such as 2-cycle detection. This fact is perhaps unsurprising because each node-adjacency tokenization input is a length-$n$ boolean vector, which must be compressed in a lossy manner if the embedding dimension $m$ and bit precision $p$ satisfy $mp=o(n)$. It is natural to ask whether such results apply to sparse graphs as well, such as graphs with bounded degree. Here, we show that requisite embedding dimension to detect 2-cycles scales linearly with the degree of the graph.
@@ 4.3 A sublinear width solution for cycle detection in bounded degree graphs
For any $n\in\mathbb{N}$ and $d\leq n$, there exists a single-layer transformer with embedding dimension $O(d\log n)$ that detects 2-cycles in any graph with node degree at most $d$. This embedding dimension is optimal up to logarithmic factors.
The proof reduces the dimensionality of the input adjacency matrix by incorporating vector embeddings from Sanford et al. (2024c) into the key, query, and value matrices to produce a “sparse attention unit,” whose activations are large when a respective cycle exists.
5 The expressive power of sub-quadratic embedding dimension
Thus far, we showed that there are graph problems that can be solved with fixed depth and linear width. A natural question is if there are problems where larger width is necessary? Intuitively, quadratic width should suffice for solving any task, since it can be used to record the entire graph, and then a sufficiently expressive MLP can solve the task. But are there problems where quadratic width is necessary. In Sec. 5.1 we show that for the problem of Eulerian cycle detection, quadratic width is necessary. This leads to the interesting question of whether there are problems that require super-linear but sub-quadratic width (i.e., between linear and quadratic). We offer a first result in this direction by showing in Sec. 5.2 that such width is sufficient for the problem of sub-graph counting.
5.1 Necessity of nearly-quadratic width for Eulerian cycle detection
As mentioned above, quadratic width trivially suffices for solving graph problems. It is therefore interesting to understand whether this upper bound is tight for some graph problems. We affirmatively answer this question with the Eulerian cycle verification problem. Given a graph and a list of “path fragments,” where each fragment consists of a pair of subsequent edges in a path, the goal of the Eulerian cycle verification problem is to determine whether the properly ordered fragments comprise an Eulerian cycle [^3] Recall that a Eulerian cycle is a cycle over the entire graph that uses each edge exactly once.. We establish the hardness of this problem on multi-graphs under the well-accepted “$1$ vs. $2$ cycle” conjecture.
@@ 5.1 Necessity of nearly-quadratic width for Eulerian cycle detection
The proof relies on a novel reduction to the $1$ vs. $2$ cycle problem. The difficulty with the reduction is that the one vs. two cycles problem consists of a sparse graph with only $n$ edges and nodes, and to show a sub-quadratic lower bound we need a dense graph with $\Omega(n^{2})$ edges and $n$ nodes. To this end, we do a random projection of an instance of the one vs. two cycles problem with $n^{2}$ nodes to a multigraph with $n$ nodes, where each node represents $n$ nodes in the original graph, but keeping all the edges. This allows us to produce a dense multigraph, for which solving the Eulerian cycle verification problem will solve the one vs. two cycles problem in the original graph, before the projection.
5.2 Sub-quadratic solutions to sub-graph counting
Subgraph detection and counting are central tasks for fields as diverse as biology, organic chemistry, and graph kernels (see Jin et al. (2020); Jiang et al. (2010); Pope et al. (2018); Shervashidze et al. (2009); see also the discussion in Chen et al. (2020)). In what follows we show that this problem can be solved with transformers with fixed depth, but width that is between linear and quadratic.
Theorem 5.2.
Let $k,n\in\mathbb{N}$, and let $G^{\prime}$ be a graph with $k$ nodes. There exists a transformer with $O(1)$ self-attention layers and embedding dimension $O\left(n^{2-1/k}\right)$ that, for any graph $G$ of size $n$, counts the number of occurrences of $G^{\prime}$ as a subgraph of $G$.
In contrast, known GNN constructions (e.g., see Chen et al., 2020) entail higher-order constructions with $k$-IGN models, which require computing practically-infeasible $k$-order tensors. These models require a depth of $k$ to recognize subgraph of size $k$, where each layer uses a $k$-tensor instead of standard matrices, which requires $n^{k}$ parameters each. Even for $k=3$ or $k=4$ this may be too large for graphs with a few hundreds of nodes.
Our proof implements the seminal “tri-tri-again” algorithm (Dolev et al., 2012) using transformers. Given a graph with $n$ nodes, we partition the nodes into $n^{1/k}$ disjoint sets, each containing $n^{1-1/k}$ nodes. For each possible combination of $k$ such sets, we use an MLP to count the given subgraph in it. There are $\binom{n^{1/k}}{k}\leq n^{k\cdot 1/k}=n$ such combinations of sets, each one containing at most $n^{2-2/k}$ edges. Thus, each token with a large enough embedding dimension can simulate one combination of subsets, cumulating all the relevant edges. Note that subgraph detection is a sub-task of counting, where the number of occurrences is larger than 0.
Theorem˜5.2 can be compared to Theorem 23 in Sanford et al. (2024a) that provides a construction for counting $k$-cliques using transformers with sub-linear memory and additional blank tokens. There, it was shown that it is possible to count $k$-cliques with a transformer of depth $O(\log\log(n))$, however the number of blank tokens is $O(n^{k-1})$ in the worse case. Here, blank tokens refer to empty tokens that are appended to the input and used for scratch space as defined in Sanford et al. (2024a). In our result, we require a depth of $O(1)$, and the total number of tokens is $n$, while the embedding dimension is super-linear (but sub-quadratic). Thus, our solution has better memory usage, since the increase in width is only polynomial, but the number of tokens is $n$ instead of $O(n^{k})$.
@@ 6 Experiments
We showcase the effect of width growth in transformers by training a variety of models on synthetic graph algorithmic tasks. Section˜6.1 establishes the relative advantages of scaling width over depth by training a family of models with similar parameter counts and variable network topologies; our results show that the shallower and wider models yield the same accuracy as the deeper models, but with faster training and inference. Section˜6.2 evaluates the critical width at which the substructure counting task becomes solvable, and shows that it is roughly linear. Section˜6.3 demonstrates the trade-offs of each graph encoding scheme (node-adjacency versus edge-list versus Laplacian-eigenvector) on different tasks. [^4] Code is provided in the Supplementary Material.
In our experiments, we used a standard transformer architecture using Pytorch’s transformer encoder layers (Paszke et al., 2019). Specifically, each layer is composed of Multi-Head Self-Attention, Feedforward Neural Network, Layer Normalization and Residual Connections. More experimental details are provided in Appendix E.
For all experiments in Section˜6.1 and Section˜6.2, we used the adjacency rows tokenization as described in Section˜5.2. Details of the implementation are described in Appendix˜D. We considered the tasks of connectivity, triangles count, and 4-cycle count. For the counting tasks, we used the substructure counting dataset from Chen et al. (2020), where each graph was labeled with the number of pre-defined substructures it contains, as a graph regression task. For the connectivity tasks, we generated synthetic graphs, and the label indicates whether the graph is connected or not. All the datasets information is described in detail in  Appendix˜E.
@@ 6.1 Empirical trade-offs between width and depth
We trained a transformer with a fixed amount of 100k parameters split between varying depth and width. We examine how the running time, loss, and generalization depend on the width and depth. We examine the following pairs of (depth, width): $(1,125),(2,89),(4,63),(8,45),(10,40)$. We train each model for $100$ epochs and examine the following: the total training time, the total inference time, the training loss and test performance (accuracy for classification and MAE for regression). We repeat this experiment with graph sizes 50 and 100. We report the averages over $3$ runs with random seeds. The hyper-parameters we tuned are provided in Appendix˜E.
The results for the connectivity task over $100$ nodes are presented in Figure˜2. Additional results for the counting tasks and different graph sizes, present the same trends, and are provided in Appendix˜D due to space limitations. As shown in Figure˜2 and Figure˜2, the training loss and accuracy remain consistent across all depth and width configurations. However, Figure˜2 reveals that shallow and wide transformers significantly reduce the total training and inference time compared to their deeper and narrower counterparts. This may be due to the ability of GPUs to parallelize the computations across the width of the same layer, but not across different layers.
6.2 Minimum width for substructure counting